OpenStudy (anonymous):

A scale drawing of a game room is shown below: A rectangle is shown. The length of the rectangle is labeled 2 inches. The width of the rectangle is labeled 4.5 inches. The scale is 1 to 30. What is the area of the actual game room in square feet? Round your answer to the nearest whole number. 9 ft2 23 ft2 56 ft2 270 ft2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@phi

OpenStudy (phi):

there are two strategies. One way: multiply both sides by 30 to get the size of the big figure. then figure out its area. 2nd way: find the area of the small figure. then multiply by 30^2 (because we know area will go up by a factor of 30*30)

OpenStudy (phi):

the other problem is that your answer will be in square inches. to change to square feet you need to divide square inches by (12*12) to get square feet.

OpenStudy (anonymous):

so the first one we get 135 and 60 tnow how do we find the area

OpenStudy (anonymous):

@phi

OpenStudy (phi):

ok, but remember that those are in inches. can you change those to feet ? (divide by 12)

OpenStudy (anonymous):

then we multiply those and get 8100 then we divide by 144 we get 56.25

OpenStudy (anonymous):

but then if we divide 135 by 12 we get 11 .25 andthen we divide 60 / 5 then we get 5 so we multiply those and get 56.2 anyway

OpenStudy (anonymous):

56.25* sorry

OpenStudy (phi):

yes, or if we first changed them to feet 135/12 * 60/12 (which by the way is \( \frac{135\cdot 60}{12\cdot 12} \) ) we would get 11.25*5= 56.25 round to 56

OpenStudy (anonymous):

ok so 56 ft^2

OpenStudy (phi):

yes, it's good (when learning) to see how to do it different ways. It might sink in !

OpenStudy (anonymous):

ok well thanks for the help today

OpenStudy (anonymous):

@iceicebaby you do flvs don't you.